Suppose 4.00 grams of H2 were used?how many grams of water would be produced?

1
Expert's answer
2022-03-24T10:09:02-0400

The reaction of water formation is:

"2H_2+O_2=2H_2O"

The mole number of hydrogen and water is equal by the chemical reaction, thus the mole number of the given mass can be used directly to compute the mass of water.

"n(H_2)=n(H_2O)"

"\frac{m(H_2)}{M(H_2)}=\frac{m(H_2O)}{M(H_2O)}"

"m(H_2O=\frac{m(H_2)\times M(H_2O)}{M(H_2)} = \frac{4.00 g \times 18 g/mol}{2 g/mol}=36.00 g"